OpenStudy (anonymous):

Determine the limit. lim as x-->0 (sinx)/(2x)

OpenStudy (abb0t):

Maybe this might help: \(\sf \frac{1}{2} \lim_{x \rightarrow 0} \frac{sin(x)}{x}\)

OpenStudy (anonymous):

well ik the answer is 1. i just dont know how

zepdrix (zepdrix):

The function sin(x) acts like x when it get's near zero.\[\Large\bf\sf \lim_{x\to0}\frac{\sin x}{2x}\quad=\quad \frac{1}{2}\lim_{x\to0}\frac{\sin x}{x}\quad\text{~}\quad \frac{1}{2}\lim_{x\to0}\frac{x}{x}\]

OpenStudy (anonymous):

the answer isn't 1, and you should use l'hopitals rule

zepdrix (zepdrix):

Pull the 1/2 out like abbot suggested, then you can maybe see what's going on. This is related to an identity that you should probably memorize if you can:\[\Large\bf\sf \lim_{x\to0}\frac{\sin x}{x}\quad=\quad 1\]

OpenStudy (anonymous):

of course the answer cannot be 1. look at the above identity and also pull out the 1/2

OpenStudy (anonymous):

The limit is x when you use radians.